# Which Retain Set Matters for LLM Unlearning? A Case Study on Entity Unlearning

### v1

@@ Abstract
Large language models (LLMs) risk retaining unauthorized or sensitive information from their training data, which raises privacy concerns. LLM unlearning seeks to mitigate these risks by selectively removing specified data while maintaining overall model performance. However, most existing work focus on methods to achieve effective forgetting and does not provide a detailed analysis of the retain set, the portion of training data that is not targeted for removal. In this paper, we investigate the effects of unlearning on various subsets of the retain set through a case study on entity unlearning. We introduce the Syntactically Similar Neighbor Set, a group of queries that share similar syntactic structures with the data targeted for removal, and show that this subset suffers the greatest performance drop during unlearning. Moreover, when used for regularization, this set not only preserves performance on syntactically similar queries but also delivers comparable or improved results across other data subsets. Our results highlight that syntactic similarity is a critical factor, potentially more so than domain or entity relationships, in achieving effective and practical LLM unlearning.
@@ 1 Introduction
To answer the research questions, we first challenge the conventional approach to neighbor set construction. While previous work Choi et al. (2024); Yuan et al. (2025) primarily focused on Domain Neighbor Sets containing instances from the same professional domain and Entity Neighbor Sets Jin et al. (2024); Choi et al. (2024) comprising closely associated entities, our research reveals that one key factor has been overlooked: Syntactic Similarity. To address this, we introduce the Syntactically Similar Neighbor Set, which contains questions sharing similar syntactic structures with the forget set. Our experiments show that this set suffers a much larger drop in performance compared to the traditional neighbor sets. (§5). This finding challenges the earlier belief Yuan et al. (2025) that entity or domain similarity is the main driver of forgetting patterns. Our paraphrasing experiments and gradient analysis confirm this result by revealing stronger interdependencies within syntactically similar information
@@ 3.3 Generating QA Pairs
For the forget set, domain neighbor set, and entity neighbor set, we employ GPT-4o to generate QA pairs for each entity. We first gather relevant passages from Wikipedia pages corresponding to each target entity. These passages serve as the context for prompting GPT-4o to generate QA pairs related to the targets. To validate the model’s knowledge and the quality of the generated pairs, we use these QA pairs to probe the evaluated model. We retain only those QA pairs for which the model successfully recalls the correct answer. This validation ensures both the consistency of the QA pairs and confirms the model’s existing knowledge.
@@ 4.3 Implementation Details
For the TOFU benchmark (Maini et al., 2024), we utilize fine-tuned Llama-2-7b-chat (Touvron et al., 2023), which has been trained on the constructed dataset to ensure it precisely answers questions in TOFU. For the real-world scenario benchmark, we employ Llama-3-8B-Instruct (Dubey et al., 2024). To enable a fair comparison of different unlearning methods at similar levels of forgetting, we adjust the hyperparameters to keep Forget Efficacy between 0.65 and 0.75. Further details are provided in Appendix C.
@@ 5.2 Exploring Domain Effects on Forgetting in Syntactically Similar Cases
To examine the domain-specific effects of unlearning in syntactically similar cases, we conduct experiments in real-world scenario across five distinct categories: Human, Company, Creative Works, Fictional Character, and Product. This analysis builds on our previous findings that syntactically similar neighbor sets exhibit more pronounced forgetting than those based on domain or entity similarity.
Prior studies Jin et al. (2024); Maini et al. (2024) define data distributions similar to $\mathcal{D}_{f}$ using either entity similarity or professional domain association, assuming that these sets would experience the most significant forgetting. Under this assumption, the Human category would be expected to exhibit the highest degree of forgetting, as entities within this category tend to be closely related. However, as shown in Figure 4, our results reveal the opposite trend—across all evaluated methods except for DPO, non-human categories exhibit substantially higher forgetting rates. This finding challenges the conventional assumption that neighbor sets constructed based on entity or domain similarity necessarily lead to higher forgetting.
@@ Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).

### v3

@@ Abstract
Large language models (LLMs) risk retaining unauthorized or sensitive information from their training data, which raises privacy concerns. LLM unlearning seeks to mitigate these risks by selectively removing specified data while maintaining overall model performance. However, most existing work focuses on methods to achieve effective forgetting and does not provide a detailed analysis of the retain set, the portion of training data that is not targeted for removal. In this paper, we investigate the effects of unlearning on various subsets of the retain set through a case study on entity unlearning. We introduce the Syntactically Similar Neighbor Set, a group of queries that share similar syntactic structures with the data targeted for removal, and show that this subset suffers the greatest performance drop during unlearning. Moreover, when used for regularization, this set not only preserves performance on syntactically similar queries but also delivers comparable or improved results across other data subsets. Our results highlight that syntactic similarity is a critical factor, potentially more so than domain or entity relationships, in achieving effective and practical LLM unlearning.
@@ 1 Introduction
To answer the research questions, we first challenge the conventional approach to neighbor set construction. While previous work Choi et al. (2024); Yuan et al. (2025) primarily focused on Domain Neighbor Sets containing instances from the same professional domain and Entity Neighbor Sets Jin et al. (2024); Choi et al. (2024) comprising closely associated entities, our research reveals that one key factor has been overlooked: Syntactic Similarity. To address this, we introduce the Syntactically Similar Neighbor Set, which contains questions sharing similar syntactic structures with the forget set. Our experiments show that this set suffers a much larger drop in performance compared to the traditional neighbor sets. (§5). This finding challenges the previous belief Yuan et al. (2025) that entity or domain similarity is the main driver of forgetting patterns. Moreover, the performance degradation is even more pronounced when syntactic similarity overlaps with entity or domain similarity, suggesting a compounding effect. Our paraphrasing experiments and gradient analysis confirm this result by revealing stronger interdependencies within syntactically similar information.
@@ 3.3 Generating QA Pairs
For the forget set, domain neighbor set, and entity neighbor set, we employ GPT-4o to generate QA pairs for each entity. We first gather relevant passages from Wikipedia pages corresponding to each target entity. These passages serve as the context for prompting GPT-4o to generate QA pairs related to the targets. Second, we further filter the QA pairs by prompting GPT-4o with the questions alone—without any passages—and retaining only those for which it produces the correct answer.
To validate the model’s knowledge and the quality of the generated pairs, we use these QA pairs to probe the evaluated model. We retain only those QA pairs for which the model successfully recalls the correct answer. This validation ensures both the consistency of the QA pairs and confirms the model’s existing knowledge.
@@ 4.3 Implementation Details
GA NPO IDK DPO
Real-world 0.734 0.745 0.657 0.721
TOFU 0.676 0.710 0.685 0.686
Table 1: Forget efficacy of each method across different scenarios.
For the TOFU benchmark (Maini et al., 2024), we utilize fine-tuned Llama-2-7b-chat (Touvron et al., 2023), which has been trained on the constructed dataset to ensure it precisely answers questions in TOFU. For the real-world scenario benchmark, we employ Llama-3-8B-Instruct (Dubey et al., 2024). To enable a fair comparison of different unlearning methods at similar levels of forgetting, we adjust the hyperparameters to keep Forget Efficacy between 0.65 and 0.75. Further details are provided in Appendix F.
(a) Real-world Scenario
(b) TOFU
@@ 5.1 Analyzing Performance Drops Across Neighbor Sets
Overlapping Sets Lead to Even Greater Forgetting. In the real-world scenario, subsets that overlap syntactic similarity with domain or entity similarity (Syn Sim & Domain, Syn Sim & Entity) experience the most severe utility drop (Figure 3(a)). This highlights that overlapping neighbor characteristics intensify forgetting effects during unlearning.
@@ 5.2 Exploring Domain Effects on Forgetting in Syntactically Similar Cases
To examine the domain-specific effects of unlearning in syntactically similar cases, we conduct experiments in real-world scenario across five distinct categories. This analysis builds on our previous findings that syntactically similar neighbor sets exhibit more pronounced forgetting than those based on domain or entity similarity.
While overlapping characteristics intensify forgetting, this raises the question of which similarity type is the primary driver. Prior studies Jin et al. (2024); Maini et al. (2024) have operated on the assumption that entity or domain similarity is the most critical factor, meaning sets with high internal similarity would be most vulnerable. Following this logic, the Human category, containing closely related entities, should exhibit the highest degree of forgetting.
However, as shown in Figure 4, the results trend in the opposite direction—non-human categories consistently exhibit substantially higher forgetting rates across most methods. This directly challenges the conventional assumption that entity or domain similarity is the most reliable predictor of performance degradation. Instead, it suggests that these factors are secondary to a more influential driver, reinforcing our central claim about the overriding importance of syntactic structure.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
@@ Appendix F Detailed Forget Quality and Model Utility for Each Method in Each Experiment
Method lr epochs Forget Efficacy EntityNeigh DomainNeigh SynNeigh
GA 5e-06 3 0.734 42.28 42.92 51.30
GA 6e-06 3 0.731 41.15 40.72 48.44
NPO 4e-05 5 0.751 47.33 43.05 52.73
NPO 3e-05 4 0.749 43.40 42.64 51.43
IDK 2e-06 4 0.650 71.63 73.45 77.14
IDK 4e-06 4 0.722 77.39 77.99 81.43
DPO 6e-06 4 0.600 48.60 47.73 57.79
DPO 7e-06 4 0.622 55.34 52.27 70.26
Table 13: Effect of hyperparameters in the real-world scenario.